OpenStudy (anonymous):

At the bank of a river, by measurement, a = 9 ft., b = 15 ft., c = 12 ft., and d = 7 ft. How long is x? How long is y?

OpenStudy (anonymous):

so a:c = d:y = b:x or 9/12 = 7/y or 12/9 = y/7 and solve as before. Also, you get 9/12 = 15/x

OpenStudy (anonymous):